OpenStudy (constanceleak):

Solve the equation for y. Solve for y 8x – 9y = 11

OpenStudy (anonymous):

11/9

OpenStudy (onepieceftw):

@johnnyy999 you're wrong cause we have two variables. We couldn't possibly find the solution.

OpenStudy (constanceleak):

actually it's a two variable solution

OpenStudy (constanceleak):

so it's possible

OpenStudy (onepieceftw):

Subtract 8x from both sides –9y = 11 - 8x Divide both sides by -9 \[y = -\frac{ 11 }{ 9 } - 8x\]

OpenStudy (constanceleak):

i thought 8x needed to be divided by 9

OpenStudy (onepieceftw):

right that's correct \[y = \frac{ -11-8x }{ -9 }\]

OpenStudy (onepieceftw):

Sorry :/

OpenStudy (constanceleak):

It's ok

OpenStudy (constanceleak):

\[y=-9/8x-9/11\]

OpenStudy (constanceleak):

y=9/8x+11/8

OpenStudy (constanceleak):

y=-8x-11

OpenStudy (constanceleak):

or y=8/9x-11/9

OpenStudy (constanceleak):

those are the only answers

OpenStudy (yoloshroom):

It's the last one :) I believe y=8/9x -11/9

OpenStudy (yoloshroom):